OpenStudy (anonymous):

what are the points of trisection of the segment joined by (-2, 8) and (10, -1)?

OpenStudy (anonymous):

its easy take the difference of 2 points then multiply them by 1/3 and 2/3

OpenStudy (anonymous):

it has to be 1/3 and 2/3 all the time?

OpenStudy (anonymous):

(10,-1)- (-2,8) = (12,9) now multiply by 1/3 12/3=4 ,9/3= 3 (4,3) and to for 2nd point multiply by 2/3 and tell me the answer

OpenStudy (anonymous):

yes that's same for all

OpenStudy (anonymous):

i figured that my working equation should something like this\[Xr= X1+r(X2-X!) & Yr=Y1+r(Y2-Y1)\] for both points

OpenStudy (anonymous):

and later add the initial points in your answers like (4,3) + (-2,8)

OpenStudy (anonymous):

oh okay then

OpenStudy (anonymous):

i think i get it now thanks